OpenStudy (anonymous):

jordan went to europe for vacation. he spent 2/3 of his time in spain. if he was in spain for 14 days, how long was hi in europe? let v represent the number of days he was europe. a.2/3v=14 b.2/3(14)=v c.v divided by 14=2/3

OpenStudy (mathstudent55):

2/3 of his vacation days is 14 days What does that suggest? The problem tells you to use v for days. Also, in math, usually "of" means times. Finally, "is" means "=".

OpenStudy (anonymous):

a ?

OpenStudy (anonymous):

or b ?

OpenStudy (amonoconnor):

It's definitely B.

OpenStudy (anonymous):

ok thank u again (:

OpenStudy (amonoconnor):

Yep!

OpenStudy (mathstudent55):

The answer is not B. 2/3 of his vacation days is 14 days 2/3 * v = 14 (2/3)v = 14 The answer is A.